OpenStudy (anonymous):

Given the function f(x) = x3 + x2 − 2x + 1, what is the resulting function when f(x) is shifted to the left 1 unit? f(x − 1) = x3 − 2x2 − x + 3 f(x + 1) = x3 + 4x2 + 3x + 1 f(x) − 1 = x3 + x2 − 2x f(x) + 1 = x3 + x2 − 2x + 2

OpenStudy (anonymous):

f(x)=(x+1)^3+ (x+1)^2-2(x+1)+1

OpenStudy (anonymous):

u can expand it and u ll get ur answer :)

OpenStudy (firekat97):

theres an easier way tho haha

OpenStudy (anonymous):

Is it C?

OpenStudy (anonymous):

cause thats what i got

OpenStudy (anonymous):

not c i ment B

OpenStudy (anonymous):

or am i wrong?

OpenStudy (firekat97):

Suppose the parabola y = x^2  thats how its graphed, now suppose y = x^2 + 1 so by adding a 1 to the function, the function is moved one spot to the left, see; f(x + 1) = (x + 1)^2 which has a zero only at x = -1 , so in the same way, by substituting (x + 1) into the above function, we can move it one spot to the left